OpenStudy (anonymous):

a company will need $65,000 in 6 years for a new addition. to meet this goal the company deposits money in an account today that pays 3% annual interest compounded quarterly. find the amount that should be invested to total $65,000 in 6 years. the company should invest..... $... round to the nearest dollar as needed.

hero (hero):

65000 = P(1+r/n)^nt

hero (hero):

65000 = P(1+.03/4)^(6*4)

OpenStudy (anonymous):

im havin a hard time solve that man

hero (hero):

65000 = P(1.0075)^(24) 65000/(1.0075)^24 = P P = 54,329

OpenStudy (anonymous):

so to round that i would just put 54,000

hero (hero):

No. Never round unless asked to

hero (hero):

That is already rounded to the nearest dollar by the way

OpenStudy (anonymous):

it says round to the nearest dollar as needed

OpenStudy (anonymous):

oh

hero (hero):

If it wasn't you would see decimals